OpenStudy (anonymous):

Thank, everyone! Got another one: 3 9 4 5 25 20 21 441 Should the next number be: 436, 442,430,439 or 454

OpenStudy (anonymous):

it should be 436. 3^2 =9 then subtract5: ( =4) the next number is 4+1 =5 5^2 = 25 and subtract 5 (=20) next no. is 21. 21^2=441 subtract 5 to get 436